Would someone please show me how to work this problem using the distribute property process. Thanks!
Two tractor trailers leave from the same point at the same time, traveling in opposite directions. One averages 52 miles per hour, and the other averages 10 miles per hour slower. In how many hours will they be 282 miles apart?

Answer by checkley75(3666)

52x+(52-10)x=282
52x+(42)x=282
52x+42x=282
94x=282
x=282/94
x=3 hours they will be 282 miles apart
proof
52*3+42*3=282
156+126=282
282=282